# thanks: Corresponding author

### v1

@@ V Conclusions
The first detection of a GW signal from a CCSN is highly anticipated as it can provide important information about the physical processes occurring during the gravitational collapse of massive stars. In this paper we have discussed a method to extract information through the analysis of simulated CCSN waveforms. In particular we have developed DL techniques to perform classification and parameter inference of rapidly rotating CCSN using the information encoded in the early (non-stochastic) part of their gravitational waveforms. To do so we have trained NNs on a catalog of numerically generated signals [20] injected into real LIGO-Virgo background noise from O2 and O3a, both in the form of spectrograms (time-frequency diagrams) and time series. The results presented in this work provide further support to the usefulness of DL methods for GW data analysis. In the specific context of CCSN, the methods discussed here seem worth of being implemented in CCSN detection (classification) and parameter estimation pipelines.
@@
We have first presented a classification test using a ResNet101 on a dataset composed of $10^{4}$ spectrograms. CCSN GW signals were injected into real noise from O2 of the LIGO Hanford, LIGO Livingston, and Virgo detectors, making up half of the dataset. The other half contained only background noise spectrograms, which had to be separated by the network’s model. The model obtained after training accomplished an accuracy of 82% and a True Positive Rate of 90%. We have also analyzed the influence of SNR and distance to the source on the network’s performance. For SNR > 30 the accuracy of the model becomes 95%. In addition, we attain more than 50% of success on the classification only for distances below 20 kpc.
@@
For the second, and main part of our study, we have considered time series. We have found that there is a big improvement in accuracy for the results in the time domain when compared with the ones with spectrograms and there is also a significant decrease in the false alarm rate. All the tests performed with time series have been done with Residual CNNs applied to datasets composed of numerically generated signals from different waveforms. The classification test was accomplished with a $10^{4}$ dataset consisting on equal parts of elements with only background noise from O3a and elements of injected signals. The network has provided a model with an accuracy of 98% with only one false positive. These results confirm the efficiency of these networks regarding classification and show that, even with a small dataset, high values of accuracy can be obtained. By further inspecting the false negatives we noticed a correlation with the SNR of the signals, as all of the misclassifications occurred for elements with SNR < 10.
@@
While the results reported in this work show the potential of Deep Learning for inferring properties of rotational CCSN they can nonetheless be improved. An obvious improvement would be accomplished by enlarging the datasets. The datasets used for the training of our models only have $10^{4}$ elements, which is considered a small size for the kind of classification and regression tests with neural networks. However, enlarging the datasets would require having available more waveforms from CCSN simulations, which are computationally costly. We note that we are currently working in the computation of new fast-rotating models for different progenitor masses and expect to present an updated study of our findings elsewhere. In addition, in this study we have used background noise from the existing network of GW detectors (namely Advanced LIGO and Advanced Virgo). As current detectors are improved and third-generation detectors become operational, the sensitivity will improve and the background noise will be reduced, which will lead to higher SNR values and easier recognition of CCSN GW signals.
@@ Acknowledgements
We thank Pablo Cerdá-Durán, Anastasios Theodoropoulos, José D. Martín, Nino Villanueva, and Pedro Passos for fruitful discussions during the course of this work. GE acknowledges the Spanish MINECO grants MINECO/FEDER Projects, PID2021-122547NB-I00 FIS2021, MCIN with funding from European Union NextGenerationEU PR47/21 MADQuantum-CM PRTR-CM and Ministry of Economic Affairs Quantum ENIA project. OGF is supported by an FCT doctoral scholarship (reference UI/BD/154358/2022). JAF and ATF are supported by the Spanish Agencia Estatal de Investigación (grant PID2021-125485NB-C21) funded by MCIN/AEI/10.13039/501100011033 and ERDF A way of making Europe. Further support is provided by the Generalitat Valenciana (grant CIPROM/2022/49), by the EU’s Horizon 2020 research and innovation (RISE) programme H2020-MSCA-RISE-2017 (FunFICO-777740), and by the European Horizon Europe staff exchange (SE) programme HORIZON-MSCA-2021-SE-01 (NewFunFiCo-101086251). AO in partially supported by FCT, under the contract CERN/FIS-PAR/0037/2021. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.

### v2

@@ V Conclusions
The first detection of a GW signal from a CCSN is highly anticipated as it can provide important information about the physical processes occurring during the gravitational collapse of massive stars. In this paper we have discussed a method to extract information through the analysis of simulated CCSN waveforms. In particular, we have developed DL techniques to perform classification and parameter inference of rapidly rotating CCSN events using the information encoded in the early (nonstochastic) part of their gravitational waveforms. To do so we have trained NNs on a catalog of numerically generated signals [20] injected into real LIGO-Virgo background noise from O2 and O3a, both in the form of spectrograms (time-frequency diagrams) and time series. The results presented in this work provide further support to the usefulness of DL methods for GW data analysis. In the specific context of CCSN, the methods discussed here seem worthy of being implemented in CCSN detection (classification) and parameter estimation pipelines.
@@
Following previous work on CBC GW signals [49], we have first presented a classification test using a ResNet101 on a dataset composed of $10^{4}$ spectrograms. CCSN GW signals were injected into real noise from O2 of the LIGO Hanford, LIGO Livingston, and Virgo detectors making up half of the dataset. All signals have a fixed distance of 20 kpc, a fixed sky position, and a duration of 200 ms. The other half of the dataset contained only background noise spectrograms, which had to be separated by the network’s model. The model obtained after training accomplished an accuracy of 82% and a precision of 90%. We have also analyzed the influence of SNR and distance to the source on the network’s performance. For SNR > 30 the accuracy of the model becomes 95%. In addition, we attain more than 50% of success on the classification only for distances below 20 kpc. While there is room for improvement, we see that using spectrograms in DL shows great promise, with results aligning with the values achieved in [49] for CBCs in O2 noise. The percentage of false positives (and false negatives) is increased due to the numerous glitches in O2 and the lower sensitivity of the detectors at that time. This issue can be minimized using background noise from more recent observing runs. As expected, when changing the background noise of the spectrograms to O3a, the accuracy of the network goes up to 95%, reflecting the better noise conditions of O3a when compared to O2.
@@
For the second and main part of our study, we have considered time series. This classification test was performed with a ResCNN on a dataset composed of $10^{4}$ time series where CCSN GW signals had been injected into 1-second-long sections of real noise from O3a of the two LIGO detectors and the Virgo detector. For this method, we have variable distance, from 5 to 20 kpc, and random sky position. All the tests performed with time series have been done with residual CNNs applied to datasets composed of numerically generated signals from different waveforms. The classification test was accomplished with a $10^{4}$ dataset consisting of equal parts of elements with only background noise from O3a and elements of injected signals. The network has provided a model with an accuracy of 98% with only one false positive. These results confirm the efficiency of these networks regarding classification and show that, even with a small dataset, high values of accuracy can be obtained. By further inspecting the false negatives we noticed a correlation with the SNR of the signals, as all of the misclassifications occurred for elements with SNR < 10. These results show the excellent performance of deep learning techniques for detection, with the time series method reflecting slightly better results than the values achieved with spectrograms in O3a.
@@
While the results reported in this work show the potential of deep learning for inferring properties of rotational CCSN events, they can nonetheless be improved. An obvious improvement would be accomplished by enlarging the datasets. The datasets used for the training of our models only have $10^{4}$ elements, which is considered a small size for the kind of classification and regression tests with neural networks. However, enlarging the datasets would require having available more waveforms from CCSN simulations, which are computationally costly. We note that we are currently working on the computation of new fast-rotating models for different progenitor masses and expect to present an updated study of our findings elsewhere. In addition, in this study, we have used background noise from the existing network of GW detectors (namely Advanced LIGO and Advanced Virgo). As current detectors are improved and third-generation detectors become operational, the sensitivity will improve and the background noise will be reduced, which will lead to higher SNR values and easier recognition of CCSN GW signals.
@@ Acknowledgements
We thank Pablo Cerdá-Durán, Anastasios Theodoropoulos, José D. Martín, Nino Villanueva, and Pedro Passos for fruitful discussions during the course of this work. G. E. acknowledges the Spanish MINECO Grants No. MINECO/FEDER, Projects No. PID2021-122547NB-I00 FIS2021, MCIN with funding from European Union NextGenerationEU PR47/21 MADQuantum-CM PRTR-CM, and Ministry of Economic Affairs Quantum ENIA Project. O. G. F. is supported by an FCT doctoral scholarship (reference UI/BD/154358/2022). J. A. F. and A. T.-F. are supported by the Spanish Agencia Estatal de Investigación (grant PID2021-125485NB-C21) funded by MCIN/AEI/10.13039/501100011033 and ERDF A way of making Europe. Further support is provided by the Generalitat Valenciana (Grant No. CIPROM/2022/49), by the EU’s Horizon 2020 research and innovation (RISE) Programme H2020-MSCA-RISE-2017 (FunFICO-777740), and by the European Horizon Europe staff exchange (SE) Programme HORIZON-MSCA-2021-SE-01 (NewFunFiCo-101086251). A. O. in partially supported by FCT, under the contract CERN/FIS-PAR/0037/2021. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.